# Deep Neural Networks: Multi-Classification and Universal Approximation

### v1

@@ Abstract.
We demonstrate that a ReLU deep neural network with a width of $2$ and a depth of $12N+4M-$ 1 layers can achieve finite sample memorization for any dataset comprising $N$ elements in $\mathbb{R}^{d}$, where $d\geq 1,$ and $M$ classes, thereby ensuring accurate classification.
By modeling the neural network as a time-discrete nonlinear dynamical system, we interpret the memorization property as a problem of simultaneous or ensemble controllability. This problem is addressed by constructing the network parameters inductively and explicitly, bypassing the need for training or solving any optimization problem.
Additionally, we establish that such a network can achieve universal approximation in $L^{p}(\Omega;\mathbb{R}_{+})$, where $\Omega$ is a bounded subset of $\mathbb{R}^{d}$ and $p\in[1,\infty)$, using a ReLU deep neural network with a width of $d+1$. We also provide depth estimates for approximating $W^{1,p}$ functions and width estimates for approximating $L^{p}(\Omega;\mathbb{R}^{m})$ for $m\geq 1$. Our proofs are constructive, offering explicit values for the biases and weights involved.
@@
Chair of Computational Mathematics, Fundación Deusto. Av. de las Universidades, 24, 48007 Bilbao, Basque Country, Spain.
@@ 1. Introduction and main results
Given a training dataset $\{x_{i},y_{i}\}_{i=1}^{N}\subset\mathcal{X}\times\mathcal{Y}$, where each $x_{i}$ represents an input data point and $y_{i}$ its corresponding label, and a model $\phi(x,\theta)$ parameterized by $\theta$, the property of finite sample memorization ([47, 48]) holds if the model $\phi$ can correctly assign the label $y_{i}$ to each training instance $x_{i}$, i.e.,
@@
We analyze the finite sample memorization property when $\phi(\cdot,\theta)$ corresponds to the output of a neural network, with $\mathcal{X}=\mathbb{R}^{d}$ where $d\geq 1$, and $\mathcal{Y}=\mathbb{R}^{m}$ where $m\geq 1$.
@@
When $\phi(x,\theta)$ is determined as the output of a continuous or discrete dynamical system, the problem can also be interpreted as an ensemble or simultaneous controllability property, ensuring that the initial data $\{x_{i}\}$ are mapped simultaneously to the corresponding targets $\{y_{i}\}$ (see [37, 32, 44, 38]).
@@
The memorization property is useful for classification or interpolation tasks involving an unknown function $f:\mathcal{X}\rightarrow\mathcal{Y}$, as it ensures point-wise fitting of $f$. Interpolation can then be employed to approximate $f$ in $L^{p}$-norms for some $p\in[1,\infty)$, in connection with universal approximation theorems, which ensure that various neural network architectures are dense in certain functional spaces [10, 13, 19, 20, 34, 35, 45, 48].
Figure 1. Classification of a two-dimensional dataset using the map $\phi$: $d=2$ and $m=1$.
@@
In this article, we present two main results. The first result shows that a ReLU deep neural network with a width of $2$ and at most $12N+4M-$ 1 layers satisfies the finite sample memorization property for $N$ input data and $M$ classes. Equivalently, the discrete dynamics generated by this deep neural network are simultaneously or ensemble controllable. This result is sharp in the sense that neural networks with a width of $1$ cannot achieve memorization.
In our proof, the neural network parameters are constructed systematically through a geometric and dynamic interpretation of the neural network’s architecture, without training or employing least squares or optimization methods. The parameters of the neural network determine hyperplanes that separate the ambient space into regions where the nonlinear activation function behaves differently in each layer. A strategic choice of the values of these parameters, iterated along the layers of the deep neural network, allows us to ensure the memorization property (see Section 2). To the best of our knowledge, the memorization capacity of this narrow but deep neural network architecture has not been previously established (see Section 1.3).
The second result demonstrates the universal approximation theorem in $L^{p}(\Omega;\mathbb{R}_{+})$, where $\Omega\subset\mathbb{R}^{d}$ is bounded and $p\in[1,\infty)$. This is achieved with a deep neural network of width $d+1$. As in our first result, we systematically construct explicit parameters of the neural network without employing any optimization procedure. Furthermore, the proof of our theorem is based on a geometrical and nonlinear construction and departs from the techniques used in previous articles that prove the universal approximation theorem with a fixed width [7, 27, 28, 33, 35]. Additionally, we provide explicit estimates for the neural network’s depth for $W^{1,p}$ functions. As a corollary, we extend our universal approximation theorem to $L^{p}(\Omega;\mathbb{R}_{+}^{m})$ and $L^{p}(\Omega;\mathbb{R}^{m})$ for $m\geq 1$, with explicit estimates for the width.
@@ 1.1. Problem formulation
$\bm{\sigma}_{j}:\mathbb{R}^{d_{j}}\rightarrow\mathbb{R}^{d_{j}},\qquad\bm{\sigma}_{j}(x)=\left(\sigma(x^{(1)})\ ,\dots,\sigma(x^{(d_{j})})\right).$
Let us take the positive integers $L,\,d,\,N,\,M$, and the dataset $\{x_{i},y_{i}\}_{i=1}^{N}\subset\mathbb{R}^{d}\times\{0,\dots,M-1\}$. We consider the following neural network architecture:
@@
where $i\in\{1,\dots,N\}$. In this context, $W_{j}\in\mathbb{R}^{d_{j}\times d_{j-1}}$ and $b_{j}\in\mathbb{R}^{d_{j}}$, for $j\in\{1,\dots,L\}$, represent the weight matrices and biases, respectively. Each $d_{j}$ determines the width of the $j$-layer, i.e., the dimension of the Euclidean space where the data reside at each iteration. The depth or the number of hidden layers of the neural network is represented by $L$, which is the number of iterations in (1.1), understood as a time-discrete dynamical system (see Figure 2). The discrete nonlinear dynamical system (1.1) is referred to as a $L-$hidden layer neural network. When $L\geq 3$, it is also known as Deep Neural Network (DNN) [5, 11].
@@
In the following, we denote the sequences of weights and biases defining the neural network (1.1) as $\mathcal{W}^{L}=\{W_{j}\}_{j=1}^{L}$ and $\mathcal{B}^{L}=\{b_{j}\}_{j=1}^{L}$, respectively. The width of the neural network is defined as $w_{max}=\max_{j\in\{1,\dots,L\}}\{d_{j}\}$, i.e., the number of neurons in the widest layer and system (1.1) is said to be a $w_{max}$-wide deep neural network. The width and depth of a neural network are determined by its architecture and serve as an intrinsic measure of its complexity and approximation capacity.
1.2. Main results.
@@
Let the integers $d,\,N,\,M\geq 1$ and consider the dataset $\{x_{i},y_{i}\}_{i=1}^{N}\subset\mathbb{R}^{d}\times\{0,\dots,M-1\}$. Assume that $x_{i}\neq x_{j}$ if $i\neq j$. Then, there exist parameters $\mathcal{W}^{L}$ and $\mathcal{B}^{L}$ with a width of $w_{max}=2$ and depth $L=2N+4M-1$ such that the input-output map of (1.1) satisfies
$\phi^{L}(\mathcal{W}^{L},\mathcal{B}^{L},x_{i})=y_{i},\qquad\text{for every }i\in\{1,\dots,N\}.$ (1.2)
Moreover, this result is sharp in the sense that the memorization property cannot be achieved with a width of $1$.
@@
  • Theorem 1.1 guarantees that there exists a $2-$wide deep neural network satisfying the finite sample memorization property, or equivalently, that system (1.1) is simultaneously controllable.
  • Obviously, Theorem 1.1 also ensures the finite sample memorization with $w_{max}\geq 2$.
@@
  • The neural network depth estimation is obtained from the constructive proof of Theorem 1.1, which is based on the worst-case scenario. This construction does not guarantee optimality in the estimated depth $L$, and for specific datasets memorization could be achieved with fewer layers.
@@
    Preconditioning of the data: Data are driven from the $d-$dimensional space to the one-dimensional one to reduce complexity.
@@
Our techniques can also be applied to other activation functions $\sigma$ under the conditions that:
@@
Note, however, that despite this generalization being possible, its practical interest is limited. The weights and biases will be more difficult to construct and will generally have larger norms compared to those required for the ReLU activation function, which ensures minimal complexity.
As a consequence of Theorem 1.1, we can analyze the case of $m-$dimensional labels for $m\geq 1$, i.e., when $\{y_{i}\}_{i=1}^{N}$ must be a subset of $M$ distinct points in $\mathbb{R}^{m}$, and get the following corollary.
@@
$\phi^{L}(\mathcal{W}^{L},\mathcal{B}^{L},x_{i})=y_{i},\qquad\text{for every }i\in\{1,\dots,N\}.$
Strategy of proof of Corollary 1.1. Corollary 1.1 directly follows from Theorem 1.1, considering $m$ deep neural networks as in Theorem 1.1, defined according to the $m$ components of the labels, and combining them in parallel to define a new neural network with vector-valued outputs (see Figure 5).
@@
In the previous results, all the labels were assumed to be non-negative. However, it is reasonable to consider the case where the signs on the labels may vary. Let $\operatorname{Im}(\phi^{L})$ denote the image of the input-output map defined by (1.1). Since $\operatorname{Im}(\phi^{L})\subset\mathbb{R}_{+}$, data cannot be mapped to negative labels by the architecture defined in (1.1). However, we can consider the richer neural network architecture
@@
where the extra parameters $A_{j}\in\mathbb{R}^{d_{j}\times d_{j}}$ for $j\in\{1,\dots,L\}$ allow the neural network, in particular, to map data to negative values. We denote by $\mathcal{A}^{L}=\{A_{j}\}_{j=1}^{L}$ the sequence of parameters.
@@
$\phi^{L}(\mathcal{A}^{L},\mathcal{W}^{L},\mathcal{B}^{L},x_{i})=y_{i},\qquad\text{for every }i\in\{1,\dots,N\}.$ (1.5)
Strategy of Proof for Corollary 1.2. The proof hinges on two key observations: first, that the architecture (1.4) can drive data to negative labels; and second, that this architecture coincides with (1.1) when $A_{j}=\text{Id}_{d_{j}}$ (the identity matrix in $\mathbb{R}^{d_{j}\times d_{j}}$). Initially, we define a set of auxiliary positive labels. Then, taking $A_{j}=\text{Id}_{d_{j}}$, we apply Theorem 1.1 to map the data points to these auxiliary labels by using $12N+4M-$ 1 layers. Finally, by using a particular matrix $A$, we construct a $2$-wide, one-layer neural network that maps the auxiliary labels to the original ones. The proof concludes by composing these two neural networks, obtaining a 2-wide neural network with $(2N+4M-1)+1$ layers.

### v2

# Constructive Universal Approximation and Finite Sample Memorization by Narrow Deep ReLU Networks

@@ Abstract.
We present a fully constructive analysis of deep ReLU neural networks for classification and function approximation tasks. First, we prove that any dataset with $N$ distinct points in $\mathbb{R}^{d}$ and $M$ output classes can be exactly classified using a multilayer perceptron (MLP) of width $2$ and depth at most $12N+4M-$ 1, with all network parameters constructed explicitly. This result is sharp with respect to width and is interpreted through the lens of simultaneous or ensemble controllability in discrete nonlinear dynamics.
Second, we show that these explicit constructions yield uniform bounds on the parameter norms and, in particular, provide upper estimates for minimizers of standard regularized training loss functionals in supervised learning. As the regularization parameter vanishes, the trained networks converge to exact classifiers with bounded norm, explaining the effectiveness of overparameterized training in the small-regularization regime.
We also prove a universal approximation theorem in $L^{p}(\Omega;\mathbb{R}_{+})$ for any bounded domain $\Omega\subset\mathbb{R}^{d}$ and $p\in[1,\infty)$, using MLPs of fixed width $d+1$. The proof is constructive, geometrically motivated, and provides explicit estimates on the network depth when the target function belongs to the Sobolev space $W^{1,p}$. We also extend the approximation and depth estimation results to $L^{p}(\Omega;\mathbb{R}^{m})$ for any $m\geq 1$.
Our results offer a unified and interpretable framework connecting controllability, expressivity, and training dynamics in deep neural networks.
@@
Chair of Computational Mathematics, University of Deusto. Av. de las Universidades, 24, 48007 Bilbao, Basque Country, Spain.
@@ 1. Introduction and main results
1.1. Motivation and summary of the main results
Given a training dataset $\{x_{i},y_{i}\}_{i=1}^{N}\subset\mathcal{X}\times\mathcal{Y}$, where each $x_{i}$ represents an input data point and $y_{i}$ its corresponding label or output, and a model $\phi(x,\theta)$ parameterized by $\theta$, the property of finite sample memorization [45, 46] holds if the model $\phi$ can correctly assign the label $y_{i}$ to each training instance $x_{i}$, i.e.,
@@ 1.1. Motivation and summary of the main results
We analyze the finite sample memorization property when $\phi(\cdot,\theta)$ corresponds to the output of a neural network, with $\mathcal{X}=\mathbb{R}^{d}$ for $d\geq 1$, and $\mathcal{Y}=\mathbb{R}^{m}$ for $m\geq 1$.
@@
When $\phi(x,\theta)$ is determined as the output of a continuous or discrete dynamical system, the problem can also be interpreted as an ensemble or simultaneous controllability problem, ensuring that the initial data $\{x_{i}\}$ are mapped simultaneously to the corresponding targets $\{y_{i}\}$ [35, 30, 42, 36].
@@
This memorization property is particularly valuable for classification and interpolation tasks involving an unknown function $f:\mathcal{X}\rightarrow\mathcal{Y}$, as it guarantees exact fitting at the data points. Once pointwise interpolation is achieved, one can extend this construction to approximate $f$ in $L^{p}$-norms for $p\in[1,\infty)$, aligning with universal approximation theorems. These results establish that various neural network architectures are dense in functional spaces, such as $L^{p}$ or Sobolev spaces, thus enabling global approximation from finite data [10, 12, 17, 18, 32, 33, 43, 46].
Figure 1. Classification of a two-dimensional dataset via the neural network map $\phi$, with input dimension $d=2$ and and scalar output $m=1$.
In this article, we present the following main results.
@@
  • The first shows one that a ReLU multilayer perceptron with a width $2$ and at most $12N+4M-$ 1 layers satisfies the finite sample memorization (or universal interpolation) property for $N$ input points and $M$ classes. Equivalently, the discrete dynamics generated by this MLP fulfils the property of simultaneous and/or ensemble controllable. This result is sharp, as memorization cannot be achieved by MLPs with width $1$.
    Our proof constructs the network parameters in a systematic manner, based on a geometric and dynamical interpretation of the neural network’s architecture. Specifically, the parameters at each layer define hyperplanes that partition the state space into regions where the nonlinear activation function induces distinct dynamical behaviors. By strategically selecting these parameters and iteratively applying them across layers, we ensure that the network satisfies the memorization property (see Section 2). To the best of our knowledge, such a constructive and purely geometric interpretation of how narrow MLPs achieve finite sample memorization has not been previously explored in the literature (see Section 1.4).
  • We then examine the implications of our constructive approach in the context of supervised training with $\ell^{2}$-regularization. Although the networks we construct are not obtained through optimization (but rather through geometric and dynamical considerations), we demonstrate that the explicit interpolating parameters from our first result can be used to establish upper bounds on the optimal value of the regularized empirical loss. In particular, we prove that minimizing a standard training objective with $\ell^{2}$-regularization produces parameter norms that are uniformly bounded by those of our explicit construction. Moreover, in the vanishing regularization limit, the resulting parameters converge to a minimal-norm interpolating network. This result offers a theoretical explanation for the behavior of trained networks in the small-$\lambda$ regime and reinforces the idea that exact data fitting can be achieved without uncontrolled growth in parameter magnitude.
  • Our final contribution establishes a universal approximation theorem in $L^{p}(\Omega;\mathbb{R}_{+})$, where $\Omega\subset\mathbb{R}^{d}$ is bounded and $p\in[1,\infty)$. This result is obtained using an MLP of fixed width $d+1$. As in our first result, the network parameters are constructed explicitly, without relying on any optimization procedure. The proof is geometrically motivated and nonlinear, marking a departure from prior approaches to universal approximation with fixed-width networks [7, 25, 26, 31, 33]. Crucially, our explicit construction allows for quantitative estimates on the required depth when the target function belongs to $W^{1,p}(\Omega;\mathbb{R}_{+})$. As a corollary, we extend this approximation result to the vector-valued settings $L^{p}(\Omega;\mathbb{R}_{+}^{m})$ and $L^{p}(\Omega;\mathbb{R}^{m})$ for any $m\geq 1$, with corresponding estimates on the necessary network width.
@@ 1.2. Problem formulation
$\bm{\sigma}_{j}:\mathbb{R}^{d_{j}}\rightarrow\mathbb{R}^{d_{j}},\qquad\bm{\sigma}_{j}(x)=\left(\sigma(x^{(1)})\ ,\dots,\sigma(x^{(d_{j})})\right).$ (1.1)
Given positive integers $L,\,d,\,N,\,M$, and the dataset $\{x_{i},y_{i}\}_{i=1}^{N}\subset\mathbb{R}^{d}\times\{0,\dots,M-1\}$. We consider the following multilayer perceptron:
@@
where $i\in\{1,\dots,N\}$. In this context, $W_{j}\in\mathbb{R}^{d_{j}\times d_{j-1}}$ and $b_{j}\in\mathbb{R}^{d_{j}}$, for $j\in\{1,\dots,L\}$, represent the weight matrices and biases, respectively. Each $d_{j}$ determines the width of the $j$-layer, i.e., the dimension of the Euclidean space where the data reside at each iteration. The depth or the number of hidden layers of the neural network is represented by $L$, which is the number of iterations in the discrete dynamical system (1.2), referred to as a $L-$hidden layer neural network (see Figure 2).
@@
In the following, we denote the sequences of weights and biases defining the neural network (1.2) as $\mathcal{W}^{L}=\{W_{j}\}_{j=1}^{L}$ and $\mathcal{B}^{L}=\{b_{j}\}_{j=1}^{L}$, respectively. The width of the neural network is defined as $w_{max}=\max_{j\in\{1,\dots,L\}}\{d_{j}\}$, i.e., the number of neurons in the widest layer. System (1.2) is said to be a $w_{max}$-wide deep neural network. The width and depth of a neural network are determined by its architecture and serve as an intrinsic measure of its complexity and approximation capacity.
1.3. Statement of the main results and strategies of the proofs.
@@
Let the integers $d,\,N,\,M\geq 1$ and consider the dataset $\{x_{i},y_{i}\}_{i=1}^{N}\subset\mathbb{R}^{d}\times\{0,\dots,M-1\}$. Assume that $x_{i}\neq x_{j}$ if $i\neq j$. Then, there exist parameters $\mathcal{W}^{L}$ and $\mathcal{B}^{L}$ with width $w_{max}=2$ and depth $L=2N+4M-1$ such that the input-output map of (1.2) satisfies
$\phi(\mathcal{W}^{L},\mathcal{B}^{L},x_{i})=y_{i},\qquad\text{forevery }i\in\{1,\dots,N\}.$ (1.3)
Moreover, this result is sharp since the memorization property cannot be achieved with width $1$.
@@
  • Theorem 1.1 guarantees that there exists a $2-$wide deep neural network satisfying the finite sample memorization property, or equivalently, that system (1.2) is simultaneously or ensemble controllable.
  • Obviously, Theorem 1.1 also ensures finite sample memorization with $w_{max}\geq 2$.
@@
  • The neural network depth estimation is obtained from the constructive proof of Theorem 1.1, which is based on the worst-case scenario. This construction does not guarantee optimality in the estimated depth $L$, and for specific datasets, memorization could be achieved with fewer layers.
@@
    Preconditioning of the data: Data is driven from the $d-$dimensional space to the one-dimensional one to reduce complexity.
@@
Let us consider the norms
$\left|\!\left|\!\left|(\mathcal{W}^{L},\mathcal{B}^{L})\right|\!\right|\!\right|_{2}:=\left(\sum_{j=1}^{L}\|W_{j}\|_{F}^{2}+\|b_{j}\|_{2}^{2}\right)^{1/2},$
and
$\left|\!\left|\!\left|(\mathcal{W}^{L},\mathcal{B}^{L})\right|\!\right|\!\right|_{\infty}:=\sup_{j\in\{1,\dots,L\}}\left\{\|W_{j}\|_{\infty},\|b_{j}\|_{\infty}\right\},$
where $\|\cdot\|_{F}$ denotes the Frobenius norm, $\|\cdot\|_{2}$ the $\ell^{2}$-norm, and $\|\cdot\|_{\infty}$ the $\ell^{\infty}$-norm. Then, due to the explicit construction of the parameters in Theorem 1.1, we obtain the following estimate for their norms.
Corollary 1.1.
Let $d,\,N,\,M>1$ be integers, and consider the dataset $\{(x_{i},y_{i})\}_{i=1}^{N}\subset B_{R_{x}}^{d}(0)\times B_{R_{y}}^{1}(0)$, where $B_{R_{x}}^{d}(0)\subset\mathbb{R}^{d}$ and $B_{R_{y}}^{1}(0)\subset\mathbb{R}$ denote balls of radius $R_{x}>0$ and $R_{y}>0$ centered at the origin. Then, the parameters $\mathcal{W}^{L}$ and $\mathcal{B}^{L}$ provided by Theorem 1.1 satisfy
$\left|\!\left|\!\left|(\mathcal{W}^{L},\mathcal{B}^{L})\right|\!\right|\!\right|_{2}$ $\leq C(1+R_{x}\sqrt{N}+R_{x}N\sqrt{M}+R_{y}M),$ (1.5)
$\left|\!\left|\!\left|(\mathcal{W}^{L},\mathcal{B}^{L})\right|\!\right|\!\right|_{\infty}$ $\leq C(R_{x}N+M+R_{y}),$ (1.6)
where $C>0$ is a constant independent of $M$, $N$, $R_{x}$, $R_{y}$, and $d$, but depends on $\min_{i\neq j\in\{1,\dots,N\}}\|x_{i}-x_{j}\|$.
Strategy of the proof of Corollary 1.1. The result follows directly from the explicit construction given in Theorem 1.1. Since the parameters $\mathcal{W}^{L}$ and $\mathcal{B}^{L}$ are specified layer by layer, the norm bounds are obtained by computing their contributions at each step and applying standard estimates.
Remark 1.2.
As mentioned in Remark 1.1, Theorem 1.1 ensures exact classification for any $w_{\max}\geq 2$ by zero-padding the parameters. Consequently, Corollary 1.1 also holds for all $w_{\max}\geq 2$.
@@
Our techniques can also be applied to other activation functions $\sigma$ that satisfy the following conditions:
@@
Note, however, that despite this generalization being possible, its practical interest is limited. The weights and biases will be more difficult to construct and will generally have larger norms compared to those provided by Corollary 1.1, which uses the ReLU activation function to ensure minimal complexity.
As a consequence of Theorem 1.1, we can address the case of $m$-dimensional labels for $m\geq 1$, that is, when $\{y_{i}\}_{i=1}^{N}$ is a subset of $M$ distinct points in $\mathbb{R}^{m}$. This leads to the following corollary.
@@
$\phi(\mathcal{W}^{L},\mathcal{B}^{L},x_{i})=y_{i},\qquad\text{forevery }i\in\{1,\dots,N\}.$
Strategy of proof of Corollary 1.2. Corollary 1.2 follows directly from Theorem 1.1 by considering $m$ independent neural networks, each constructed as in Theorem 1.1 to handle one component of the $m$-dimensional labels, and combining them in parallel to obtain a single network with vector-valued outputs (see Figure 5).
@@
In the previous results, all the labels were assumed to be non-negative. However, it is reasonable to consider the case where the signs on the labels may vary. Let $\operatorname{Im}(\phi^{L})$ denote the image of the input-output map defined by (1.2). Since $\operatorname{Im}(\phi^{L})\subset\mathbb{R}_{+}$, data cannot be mapped to negative labels by the architecture defined in (1.2). However, we can consider the following MLP
@@
where the extra parameters $A_{j}\in\mathbb{R}^{d_{j}\times d_{j}}$ for $j\in\{1,\dots,L\}$ allow the neural network, in particular, to map data to negative values. We denote by $\mathcal{A}^{L}=\{A_{j}\}_{j=1}^{L}$ this new sequence of parameters.
@@
$\phi(\mathcal{A}^{L},\mathcal{W}^{L},\mathcal{B}^{L},x_{i})=y_{i},\qquad\text{for every }i\in\{1,\dots,N\}.$ (1.8)
Strategy of Proof for Corollary 1.3. The proof hinges on two key observations: first, that the architecture (1.7) can drive data to negative labels; and second, that this architecture coincides with (1.2) when $A_{j}=\text{Id}_{d_{j}}$ (the identity matrix in $\mathbb{R}^{d_{j}\times d_{j}}$). Initially, we define a set of auxiliary positive labels. Then, taking $A_{j}=\text{Id}_{d_{j}}$, we apply Theorem 1.1 to map the data points to these auxiliary labels by using $12N+4M-$ 1 layers. Finally, by using a particular matrix $A$, we construct a $2$-wide, one-layer neural network that maps the auxiliary labels to the original ones. The proof concludes by composing these two neural networks, obtaining a 2-wide neural network with $2N+4M$ layers.